38) Assume the following inverted demand function of a firm in the short run: P = 50 - .5Q. Now assume the total cost function of this firm is : TC = 50 + 100Q - Q2 The above cost function yields the MC function as 100- 2Q (a). Calculate the profit maximizing price and output of this firm. (Hint: Obtain the MR first). (b). Is this firm earning a profit or loss in the short run? Explain. Is this firm earning a profit or incurring a loss? What is the amount of short-run profit or loss? Explain fully

Please type out answers for each question, because I can't read everyone handwriting please.
Answer #1

a.

P=50-0.5Q

TR=P*Q

=(50-0.5Q)Q

=50Q-0.5Q2

Taking derivative of TR with respect to Q

dTR/dQ=50(1)-1Q

MR=50-Q

MC= 100- 2Q

Profit-maximizing condition is

MR=MC

50-Q=100-2Q

2Q-Q=100-50

Q=50 units.

Substituting Q into price equation

P=50-0.5(50)

=50-25

=$25

b.

TR=P*Q

=25*50

=1,250

TC = 50 + 100Q - Q2

=50+100*50-502

=50+5000-2500

TC=2550

Since TC is greater than TR in short-run. Hence there is loss.

Loss=TC-TR

=2550-1250

=$1300